40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 51 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  14554   Fri Apr 19 11:36:23 2019 gautamUpdateSUSNo consistent solution for output matrix

Ther isn't a consistent set of OSEM coil gains that explains the best actuation vectors we determined yesterday. Here are the explicit matrices:

  1. POS (tuned to minimize excitation at ~13.5 Hz in the Oplev PIT and YAW error signals): \begin{bmatrix} \text{UL} & \text{UR} & \text{LL} & \text{LR} \end{bmatrix}\begin{bmatrix} 0.98 \\ 0.96 \\ 1.04 \\ 1.02 \\ \end{bmatrix}
  2. PIT (tuned to minimize cross coupled peak in the Oplev YAW error signal at ~10.5 Hz): ​\begin{bmatrix} \text{UL} & \text{UR} & \text{LL} & \text{LR} \end{bmatrix}\begin{bmatrix} 0.64 \\ 1.12 \\ -1.12 \\ -0.64 \\ \end{bmatrix}
  3. YAW (tuned to minimize cross coupled peak in the Oplev PIT error signal at ~13.5 Hz): \begin{bmatrix} \text{UL} & \text{UR} & \text{LL} & \text{LR} \end{bmatrix}\begin{bmatrix} 1.5 \\ -0.5 \\ 0.5 \\ -1.5 \\ \end{bmatrix}

There isn't a solution to the matrix equation \begin{bmatrix} \alpha_1 & \alpha_2 & \alpha_3 & \alpha_4 \end{bmatrix} \begin{bmatrix} 1 & 1 & 1 \\ 1 & 1 & -1 \\ 1 & -1 & 1 \\ 1 & -1 & -1 \end{bmatrix} =\begin{bmatrix} 0.98 & 0.64 & 1.5 \\ 0.96 & 1.12 & -0.5 \\ 1.04 & -1.12 & 0.5 \\ 1.02 & -0.64 & -1.5 \end{bmatrix}, i.e. we cannot simply redistribute the actuation vectors we found as gains to the coils and preserve the naive actuation matrix. What this means is that in the OSEM coil basis, the actuation eigenvectors aren't the naive ones we would think for PIT and YAW and POS. Instead, we can put these custom eigenvectors into the output matrix, but I'm struggling to think of what the physical implication is. I.e. what does it mean for the actuation vectors for PIT, YAW and POS to not only be scaled, but also non-orthogonal (but still linearly independent) at ~10 Hz, which is well above the resonant frequencies of the pendulum? The PIT and YAW eigenvectors are the least orthogonal, with the angle between them ~40 degrees rather than the expected 90 degrees.

Quote:

So we now have matrices that minimize the cross coupling between these DoFs - the idea is to back out the actuation coefficients for the 4 OSEM coils that gives us the most diagonal actuation, at least at AC. 

  14553   Fri Apr 19 09:42:18 2019 KojiBureaucracyGeneralItem borrowing (40m->OMC)

Apr 16, 2019
Borrowed two laser goggles from the 40m. (Returned Apr 29, 2019)
Apr 19, 2019
Borrowed from the 40m:
- Universal camera mount
- 50mm CCD lens
- zoom CCD lens (Returned Apr 29, 2019)
- Olympus SP-570UZ (Returned Apr 29, 2019)
- Special Olympus USB Cable (Returned Apr 29, 2019)

 

  14552   Thu Apr 18 23:10:12 2019 gautamUpdateLoss MeasurementX arm misaligned

Yehonathan wanted to take some measurements for loss determination. I misaligned the X arm completely and we installed a PD on the AS table so there is no light reaching the AS55 and AS110 PDs. Yehonathan will post the detailed elog.

  14551   Thu Apr 18 22:35:23 2019 gautamUpdateSUSETMY actuator diagnosis

[rana, gautam]

Rana did a checkout of my story about oddness of the ETMY suspension. Today, we focused on the actuators - the goal was to find the correct coefficients on the 4 face coils that would result in diagonal actuation (i.e. if we actuate on PIT, it only truly moves the PIT DoF, as witnessed by the Oplev, and so on for the other DoFs). Here are the details:

  1. Ramp times for filter modules:
    • All the filter modules in the output matrix did not have ramp times set.
    • We used python, cdsutils and ezca to script the writing of a 3 second ramp to all the elements of the 5x6 output matrix.
    • The script lives at /opt/rtcds/caltech/c1/scripts/cds/addRampTimes.py, can be used to implement similar scripts to initialize large numbers of channels (limiters, ramp times etc).
  2. Bounce mode checkout:
    • ​The motivation here was to check if there is anomalously large coupling of the bounce mode to any of the other DoFs for ETMY relative to the other optics
    • The ITMs have a different (~15.9 Hz) bounce mode frequency compared to the ETMs (~16.2 Hz).
    • I hypothesize that this is because the ETMs were re-suspended in 2016 using new suspension wire.
    • We should check out specs of the wires, look for either thickness differences or alloying composition variation (Steve has already documented some of this in the elog linked above). Possibly also check out the bounce mode for a 250g load on the table top.
  3. Step responses for PIT and YAW
    • With the Oplevs disabled (but other local damping loops engaged), we applied a step of 100 DAC counts to the PIT and YAW DoFs from the realtime system (one at a time)
    • We saw significant cross-coupling of the YAW step coupling to PIT, at the level of 50%.
  4. OSEM coil coefficient balancing
    • I had done this a couple of months ago looking at the DC gain of the 1/f^2 pendulum response.
    • Rana suggested an alternate methodology 
      • we used the lock-in amplifier infrastructure on the SUS screens to drive a sine wave
      • Frequencies were chosen to be ~10.5 Hz and ~13.5 Hz, to be outside the Oplev loop bandwidth
      • Tests were done with the Oplev loop engaged. The Oplev error signal was used as a diagnostic to investigate the PIT/YAW cross coupling.
      • In the initial tests, we saw coupling at the 20% level. If the Oplev head is rotated by 0.05 rad relative to the "true" horizontal-vertical coordinate system, we'd expect 5% cross coupling. So this was already a red flag (i.e. it is hard to believe that Oplev QPD shenanigans are responsible for our observations). We decided to re-diagonalize the actuation.
      • The output matrix elements for the lock-in-amplifier oscillator signals were adjusted by adding some amount of YAW to the PIT elements (script lives at /opt/rtcds/caltech/c1/scripts/SUS/stepOutMat.py), and vice versa, and we tried to reduce the height of the cross-coupled peaks (viewed on DTT using exponential weighting, 4 avgs, 0.1 Hz BW - note that the DTT cursor menu has a peak find option!). DTT Template saved at /users/Templates/SUS/ETMY-actDiag.xml
      • This worked really well for minimizing PIT response while driving YAW, not as well for minimizing YAW in PIT. 
      • Next, we added some YAW to a POS drive to minimize the any signal at this drive frequency in the Oplev YAW error signal. Once that was done, we minimized the peak in the Oplev PIT error signal by adding some amount of PIT actuation.
      • So we now have matrices that minimize the cross coupling between these DoFs - the idea is to back out the actuation coefficients for the 4 OSEM coils that gives us the most diagonal actuation, at least at AC. 
  5. Next steps:
    • All of our tests tonight were at AC - once the coil balancing has been done at AC, we have to check the cross coupling at DC. If everything is working correctly, the response should also be fairly well decoupled at DC, but if not, we have to come up with a hypothesis as to why the AC and DC responses are different.
    • Can we gain any additional info from driving the pringle mode and minimizing it in the Oplev error signals? Or is the problem overconstrained?
    • After the output matrix diagonalization is done, drive the optic in POS, PIT and YAW, and construct the input matrix this way (i.e. transfer function), as an alternative to the usual free-swinging ringdown method. Look at what kind of an input matrix we get.
    • Repeat the free-swinging ringdown with the ETMY bias voltage adjusted such that all the OSEM PDmons report ~100 um different position from the "nominal" position (i.e. when the Y arm cavity is aligned). Investigate whether the resulting eigenmode frequencies / Qs are radically different. I'm setting the optic free-swinging on my way out tonight. Optic kicked at 1239690286.
  14550   Wed Apr 17 18:12:06 2019 gautamUpdateVACVac interlock tripped again

After getting the go ahead from Chub and Jon, I restored the Vacuum state to "Vacuum normal", see Attachment #1. Steps:

  1. Interlock code modifications
    • Backed up /opt/target/python/interlocks/interlock_conditions.yaml to /opt/target/python/interlocks/interlock_conditions_UPS.yaml
    • The "power_loss" condition was removed for every valve and pump inside /opt/target/python/interlocks/interlock_conditions.yaml
    • The interlock service was restarted using sudo systemctl restart interlock.service
    • Looking at the status of the service, I saw that it was dying ~ every 1 second.
    • Traced this down to a problem in/opt/target/python/interlocks/interlock_conditions.yaml  when the "pump_managers" are initialized - the way this is coded up, doesn't play nice if there are no conditions specified in the yaml file. For now, I just commented this part out. The git diff  below:
  2. Restoring vacuum normal:
    • Spun up TP1, TP2 and TP3
    • Opened up foreline of TP1 to TP2, and then opened main volume to TP1
    • Opened up annulus foreline to TP3, and then opened the individual annular volumes to TP3.
controls@c1vac:/opt/target/python/interlocks$ git diff interlock.py
diff --git a/python/interlocks/interlock.py b/python/interlocks/interlock.py
index 28d3366..46a39fc 100755
--- a/python/interlocks/interlock.py
+++ b/python/interlocks/interlock.py
@@ -52,8 +52,8 @@ class Interlock(object):
         self.pumps = []
         for pump in interlocks['pumps']:
             pm = PumpManager(pump['name'])
-            for condition in pump['conditions']:
-                pm.register_condition(*condition)
+            #for condition in pump['conditions']:
+            #    pm.register_condition(*condition)
             self.pumps.append(pm)

So far the pressure is coming down smoothly, see Attachment #2. I'll keep an eye on it.

PSL shutter was opened at 645pm local time. IMC locked almost immediately.

Update 11pm: The pressure has reached 8.5e-6 torr without hiccup. 

Attachment 1: Screenshot_from_2019-04-17_18-11-45.png
Screenshot_from_2019-04-17_18-11-45.png
Attachment 2: Screenshot_from_2019-04-17_18-21-30.png
Screenshot_from_2019-04-17_18-21-30.png
  14549   Wed Apr 17 11:01:49 2019 gautamUpdateALSLarge 2kHz peak (and harmonics) in ALS X no more

EX green stayed locked to XARM length overnight without a problem. The spectrogram doesn't show any alarming time varying features around 2 kHz (or at any other frequency).

Attachment 1: EX_PDH_specGram.pdf
EX_PDH_specGram.pdf
  14548   Wed Apr 17 00:50:17 2019 gautamUpdateALSLarge 2kHz peak (and harmonics) in ALS X no more

I looked into this issue today. Initially, my thinking was that I'd somehow caused clipping in the beampath somewhere which was causing this 2kHz excitation. However, on looking at the spectrum of the in-loop error signal today (Attachment #1), I found no evidence of the peak anymore!

Since the vacuum system is in a non-nominal state, and also because my IR ALS beat setup has been hijacked for the MZ interferometer, I don't have an ALS spectrum, but the next step is to try single arm locking using the ALS error signal. To investigate whether the 2kHz peak is a time-dependent feature, I left the EX green locked to the arm (with the SLOW temperature offloading servo ON), hopefully it stays locked overnight...

Quote:

These weren't present last week. The peaks are present in the EX PDH error monitor signal, and so are presumably connected with the green locking system. My goal tonight was to see if the arm length control could be done using the ALS error signal as opposed to POX, but I was not successful.

Attachment 1: EX_PDHnoise.pdf
EX_PDHnoise.pdf
  14547   Wed Apr 17 00:43:38 2019 gautamUpdateFrequency noise measurementMZ interferometer ---> DAQ
  1. Delay fiber was replaced with 5m (~30 nsec delay)
    • The fringing of the MZ was way too large even with the free running NPRO (~3 fringes / sec)
    • Since the V/Hz is proportional to the delay, I borrowed a 5m patch cable from Andrew/ATF lab, wrapped it around a spool, and hooked it up to the setup
    • Much more satisfactory fringing rate (~1 wrap every 20 sec) was observed with no control to the NPRO
  2. MZ readout PDs hooked up to ALS channels
    • To facilitate further quantitative study, I hooked up the two PDs monitoring the two ports of the MZ to the channels normally used for ALS X.
    • ZHL3-A amps inputs were disconnected and were turned off. Then cables to their outputs were highjacked to pipe the DC PD signals to the 1Y3 rack
    • Unfortunately there isn't a DQ-ed fast version of this data (would require a model restart of c1lsc which can be tricky), but we can already infer the low freq fringing rate from overnight EPICS data and also use short segments of 16k data downloaded "live" for the frequency noise measurement.
    • Channels are C1:ALS-BEATX_FINE_I_IN1 and C1:ALS-BEATX_FINE_Q_IN1 for 16k data, and C1:ALS-BEATX_FINE_I_INMON and C1:ALS-BEATX_FINE_I_INMON for 16 Hz.

At some point I'd like to reclaim this setup for ALS, but meantime, Anjali can work on characterization/noise budgeting. Since we have some CDS signals, we can even think of temperature control of the NPRO using pythonPID to keep the fringe in the linear regime for an extended period of time.

  14546   Tue Apr 16 22:06:51 2019 gautamUpdateVACVac interlock tripped again

This happened again, about 30,000 seconds (~2:06pm local time according to the logfile) ago. The cited error was the same -

2019-04-16 14:06:05,538 - C1:Vac-error_status => VA6 closed. AC power loss.

Hard to believe there was any real power loss, nothing else in the lab seems to have been affected so I am inclined to suspect a buggy UPS communication channel. The PSL shutter was not closed - I believe the condition is for P1a to exceed 3 mtorr (it is at 1 mtorr right now), but perhaps this should be modified to close the PSL shutter in the event of any interlock tripping. Also, probably not a bad idea to send an email alert to the lab mailing list in the event of a vac interlock failure.

For tonight, I only plan to work with the EX ALS system anyways so I'm closing the PSL shutter, I'll work with Chub to restore the vacuum if he deems it okay tomorrow.

Attachment 1: Screenshot_from_2019-04-16_22-05-47.png
Screenshot_from_2019-04-16_22-05-47.png
Attachment 2: Screenshot_from_2019-04-16_22-06-02.png
Screenshot_from_2019-04-16_22-06-02.png
  14545   Mon Apr 15 22:55:34 2019 gautamFrogsThermal CompensationLab thermostat adjusted

It is feeling cold in the office area. According to the digital wall clock near the coffee machine, it is 19C. Rana bumped the thermostat setpoint up by 2F (from 75F to 77F). We need to setup long-term monitoring.

  14544   Mon Apr 15 22:39:10 2019 gautamUpdateFrequency noise measurementAlternate setup with PSL pickoff

[anjali, gautam]

just main points, anajli is going to fill out the details.

To rule out mode-matching as the reason for non-ideal output from the MZ, I suggested using the setup I have on the NW side of the PSL enclosure for the measurement. This uses two identical fiber collimators, and the distance between collimator and recombination BS is approximately the same, so the spatial modes should be pretty well matched. 

The spooled fiber we found was not suitable for use as it had a wide key connector and I couldn't find any wide-key FC/PC to narrow-key FC/APC adaptors. So we decided to give the fiber going to the Y end and back (~90m estimated length) a shot. We connected the two fibers at the EY table using a fiber mating sleeve (so the fiber usually bringing the IR pickoff from EY to the PSL table was disconnected from its collimator). 

In summary, we cannot explain why the contrast of the MZ is <5%. Spatial mode-overlap is definitely not to blame. Power asymmetry in the two arms of the MZ is one possible explanation, could also be unstable polarization, even though we think the entire fiber chain is PM. Anjali is investigating.

 


We saw today that the Thorlabs PM beam splitters (borrowed from Andrew until our AFW components arrive) do not treat the two special axes (fast and slow) of the fiber on equal footing. When we coupled light into the fast axis, we saw huge asymmetry between the two split arms of the beamsplitter (3:1 ratio in power instead of the expected 1:1 for a 50/50 BS). Looking at the patch cord with an IR viewer, we also saw light leaking through the core along it. Turns out this part is meant to be used with light coupled to the slow axis only.

  14543   Mon Apr 15 18:29:07 2019 ranaSummaryComputersnew laptop setup: ASIA - yum issues

had trouble using YUM to update. This turned out to be a config problem with our Martian router, not the new laptop. Since I've changed the WiFi pwd awhile ago for the martian access for the CDS laptops, you'll have to enter that in order to use the laptops.

turned out to be some Access Control nonsense inside of the router. Even loggin in as admin with a cable gave some of the fields the greyed out color (had to hover over the link and then type the URL directly in the browser window). ASIA is now able to connect and use YUM + usual connections. Gautam and I have also moved the router a little to get easier view of its LED lights and not blockk its WiFi signal with the cable tray. We'll get a little shelf so that we can mount it ~1 foot off of the wall.

still, this seems like a bad laptop choice: the Lenovo Ideapad 330 will not have its touchpad supported by SL7 without compiling a new version of the kernel frown

  14542   Mon Apr 15 18:13:23 2019 ranaUpdateComputer Scripts / ProgramsAutomated suspension testing with susPython

if this gonna be general for everybody, maybe it oughta be called IFOtest (like the last incarnation that was tried in Livingston) ?

:

In anticipation of needing to test hundreds of suspension signals after the c1susaux upgrade, I've started developing a Python package to automate these tests: susPython

 

then the SUS tests could just be some smaller set of measurements. Using the same code base, but different params.

  14541   Mon Apr 15 10:20:44 2019 gautamUpdateOptical LeversBS Oplev PIT was oscillating

The AS spot on the camera was oscillating at ~3 Hz. Looking at the Oplevs, the culprit was the BS PIT DoF. Started about 12 hours ago, not sure what triggered it. I disabled Oplev damping, and waited for the angular motion to settle down a bit, and then re-enabled the servo - damps fine now...

Attachment 1: BS_OL_oscillating.png
BS_OL_oscillating.png
  14540   Fri Apr 12 01:22:27 2019 AnjaliUpdateFrequency noise measurementFrequency noise measurement of 1 micron source

The alignement was disturbed after the replcement of the beam splitter. We tried to get the alignment back . But we are not succeeded yet in getting good interfernce pattern. This is mainly because of poor mode matching of two beams. We will also try with the spooled fiber.

Quote:

 

  • The experimental setup is then modified as shown in attachment #3. The thick beam spliiter is replaced with a thinner one. The mount is also changed such that the transmitted beam can be now coupled to an other photodiode (earlier  the transmitted light was blocked by the mount). One more photodiode (PDA55) is introduced .So now the two photodiodes in the setup are PDA520 and PDA 55. 
 
 
  14539   Thu Apr 11 17:30:45 2019 JonUpdateSUSAutomated suspension testing with susPython

Summary

In anticipation of needing to test hundreds of suspension signals after the c1susaux upgrade, I've started developing a Python package to automate these tests: susPython

https://git.ligo.org/40m/suspython

The core of this package is not any particular test, but a general framework within which any scripted test can be "nested." Built into this framework is extensive signal trapping and exception handling, allowing actuation tests to be performed safely. Namely it protects against crashes of the test scripts that would otherwise leave the suspensions in an arbitrary state (e.g., might destroy alignment).

Usage

The package is designed to be used as a standalone from the command line. From within the root directory, it is executed with a single positional argument specifying the suspension to test:

$ python -m suspython ITMY

Currently the package requires Python 2 due to its dependence on the cdsutils package, which does not yet exist for Python 3.

Scripted Tests

So far I've implemented a cross-consistency test between the DC-bias outputs to the coils and the shadow sensor readbacks. The suspension is actuated in pitch, then in yaw, and the changes in PDMon signals are measured. The expected sign of the change in each coil's PDMon is inferred from the output filter matrix coefficients. I believe this test is sensitive to two types of signal-routing errors: no change in PDMon response (actuator is not connected), incorrect sign in either pitch or yaw response, or in both (two actuators are cross-wired).

The next test I plan to implement is a test of the slow system using the fast system. My idea is to inject a 3-8 Hz excitation into the coil output filter modules (either bandlimited noise or a sine wave), with all coil outputs initially disabled. One coil at a time will be enabled and the change in all VMon signals monitored, to verify the correct coil readback senses the excitation. In this way, a signal injected from the independent and unchanged fast system provides an absolute reference for the slow system.

I'm also aware of ideas for more advanced tests, which go beyond testing the basic signal routing. These too can be added over time within the susPython framework.

  14538   Thu Apr 11 12:57:48 2019 JonUpdateSUSStarting some scripted SUS tests on ITMY

Testing is finished.

Quote:

Will advise when I'm finished, will be by 1 pm for ALS work to begin.

  14537   Thu Apr 11 12:21:01 2019 not KojiUpdatePSLPSL fan is noisy

I could probably install the new fan if we have one.  Can you do without the laser for a while?

Quote:

This thread: ELOG 10295

My interpretation of these ELOGs is that we did not have the replacement, and then I brought unknown fan from WB. At the same time, Steve ordered replacement fans which we found in the blue tower yesterday.
The next action is to replace the internal fan, I believe.

  14536   Thu Apr 11 12:04:43 2019 JonUpdateSUSStarting some scripted SUS tests on ITMY

Will advise when I'm finished, will be by 1 pm for ALS work to begin.

  14535   Thu Apr 11 11:42:10 2019 KojiUpdatePSLPSL fan is noisy

This thread: ELOG 10295

My interpretation of these ELOGs is that we did not have the replacement, and then I brought unknown fan from WB. At the same time, Steve ordered replacement fans which we found in the blue tower yesterday.
The next action is to replace the internal fan, I believe.

  14534   Thu Apr 11 09:05:06 2019 AnjaliUpdateIOOSpooled fiber
  • Attchment #1,2,3 and 4 shows the results with frequency modulation of 32 Hz, 140 Hz , 300 Hz and without frequency modulation. I am trying to understand these results better.
  • A lot of fringing is there even when no modulation is applied. We hope to improve this by spooling the fiber and then encasing it in a box. 
  • As mentioned by Gautam, we have got a 50 m spooled fiber. Attachment #5 shows the photo of the same
Quote:

Steve had showed me some stock of long fibers a while back - they are from Oz Optics, and are 50m long, and are already spooled - so barring objections, we will try the MZ setup with the spooled fiber and see if there is any improvement in the fringing rate of the MZ. Then we can evaluate what additional stabilization of the fiber length is required. Anjali will upload a photo of the spooled fiber.

Attachment 1: Frequecy_modulation_32_Hz.pdf
Frequecy_modulation_32_Hz.pdf
Attachment 2: Frequecy_modulation_140_Hz.pdf
Frequecy_modulation_140_Hz.pdf
Attachment 3: Frequecy_modulation_300_Hz.pdf
Frequecy_modulation_300_Hz.pdf
Attachment 4: Without_modulation.pdf
Without_modulation.pdf
Attachment 5: New_fiber_spool.JPG
New_fiber_spool.JPG
  14533   Thu Apr 11 01:10:05 2019 gautamUpdateALSLarge 2kHz peak (and harmonics) in ALS X

These weren't present last week. The peaks are present in the EX PDH error monitor signal, and so are presumably connected with the green locking system. My goal tonight was to see if the arm length control could be done using the ALS error signal as opposed to POX, but I was not successful.

Attachment 1: EX_PDH_2kNoise.pdf
EX_PDH_2kNoise.pdf
  14532   Wed Apr 10 23:37:59 2019 gautamUpdatePSLPSL fan is noisy

Attached is my phone recording of what it sounds like right now in the PSL enclosure - not good for frequency noise measurement! The culprit is the little PC fan that is hacked onto the back of the Innolight controller. 

  1. Is this necessary?
  2. If so, is it sufficient to replace this fan with one from our stock?
Attachment 1: New_Recording.m4a
  14531   Wed Apr 10 22:59:22 2019 gautamUpdateIOOSpooled fiber

Steve had showed me some stock of long fibers a while back - they are from Oz Optics, and are 50m long, and are already spooled - so barring objections, we will try the MZ setup with the spooled fiber and see if there is any improvement in the fringing rate of the MZ. Then we can evaluate what additional stabilization of the fiber length is required. Anjali will upload a photo of the spooled fiber.

  14530   Wed Apr 10 16:58:54 2019 ranaUpdateIOOfiber MZ for NPRO freq noise measurements

Can we get some panel mount FC/APC connectors and put them on a box?   Then we could have the whole setup inside of a box that is filled with foam and sits outside the PSL hut. cheeky

  14529   Wed Apr 10 00:33:09 2019 AnjaliUpdateFrequency noise measurementFrequency noise measurement of 1 micron source
  • Attachement #1 shows the input (ch4-green) modulation frequency and the photodiode output (ch1-yellow) when the modulation frequency is about 100 Hz
  • Attachement #2 shows the input (ch4-green) modulation frequency and the photodiode output (ch1-yellow) when the modulation frequency is about 30 Hz
  • The output frequency is varying in accordance with variation in modulation frequency. It is observed that, for a given modulation frequency also, the output frequency is fluctuating. There could be multiple reasons for this behaviour. One of the main reasons is the frequency noise of the laser itself. Also, there could be acoustic noise coupled to the system (eg, by change in length of the fiber).
  • The experimental setup is then modified as shown in attachment #3. The thick beam spliiter is replaced with a thinner one. The mount is also changed such that the transmitted beam can be now coupled to an other photodiode (earlier  the transmitted light was blocked by the mount). One more photodiode (PDA55) is introduced .So now the two photodiodes in the setup are PDA520 and PDA 55. 
  • We then applied frequency modulation on the input laser and observed the output of the two photodiodes. But we didn't get the results as we expected and observed earlier (shown in attachment #1 &2). Looks like, the problem is poor mode matching between the two beams. 
Quote:
  • The alignment of the output beam from the delayed path of MZI to the photodetector was disturbed when we did the polarisation characterisation yesterday. So, today we tried to align the output beam from the delayed path of MZI to the detector .
  • We then observed the beat output from the detector on oscilloscope.We initialy observed a dc shift . We then applied a frequency modulation on the input laser and observed the output on oscilloscope. We expected to see variation in output frequency in accordance with variation of input frequency modulation. But we didnt observe this and we were not really getting the interference pattern. 
  • We tried to make the alignment better. With a better alignment, we could see the interference pattern. We also observed that the output frequency was varying in accordance with variation in the input frequency modulation. We would expect a better result with proper mode matching of the two beams on the photodetector.
Attachment 1: Modulation_frequency_100Hz.jpg
Modulation_frequency_100Hz.jpg
Attachment 2: Modulation_frequency_30Hz.jpg
Modulation_frequency_30Hz.jpg
Attachment 3: Modified_setup.JPG
Modified_setup.JPG
  14528   Tue Apr 9 19:07:12 2019 gautamUpdateALSIR ALS noise budget

Updated the noise budget to include the unsuppressed frequency noise from the EX laser. It does not explain the noise between 10-100 Hz, although the 1-3 Hz noise is close.

Actually, I think the curve that should go on the budget is when the X arm length is locked to the PSL frequency, whereas this is when the X arm is just locally damped. I will update it later tonight.

Update 1010pm: I've uploaded the relevant plot as Attachment #2. Predictably, the unsuppressed frequency noise of the EX laser is now higher, because the MC length is a noisier frequency reference than the arm cavity. But still it is a factor of 10 below the measured ALS noise.

Quote:

Next noises to budget:

  1. In-loop X arm length noise
  2. In-lop EX laser frequency noise
Attachment 1: ALS_noiseBudget.pdf
ALS_noiseBudget.pdf
Attachment 2: ALS_noiseBudget.pdf
ALS_noiseBudget.pdf
  14527   Tue Apr 9 18:44:00 2019 gautamUpdateALSEX Green PDH discriminant measurement

I decided to use the more direct method, of disconnecting feedback to the EX laser PZT, and then looking at the cavity flashes. 

Attachment #1 shows the cavity swinging through two resonances (data collected via oscilloscope). Traces are for the demodulated PDH error signal (top) and the direct photodiode signal (bottom). The traces don't look very clean - I wonder if some saturation / slew rate effects are at play, because we are operating the PD in the 30 dB setting, where the bandwidth of the PD is spec-ed as 260 kHz, whereas the dominant frequency component of the light on the PD is 430 kHz.

The asymmetric horns corresponding to the sideband resonances were also puzzling. Doing the modeling, Attachment #2, I think this is due to the fact that the demodulation phase is poorly set. The PDH modulation frequency is only ~5x the cavity linewidth, so both the real and imaginary parts of the cavity reflectivity contribute to the error signal. If this calculation is correct, we can benefit (i.e. get a larger PDH discriminant) by changing the demod phase by 60 degrees. However, for 230 kHz, it is impractical to do this by just increasing cable length between the function generator and mixer.

Anyway, assuming that we are at the phi=30 degree situation (since the measurement shows all 3 horns going through roughly the same voltage swing), the PDH discriminant is ~40 uV/Hz. In lock, I estimate that there is ~60 uW of light incident on the PDH reflection photodiode. Using the PD response of 0.2 A/W, transimpedance of 47.5 kohm, and mixer conversion loss of 6dB, the shot-noise limited sensitivity is 0.5 mHz/rtHz. The photodiode dark noise contribution is a little lower - estimated to be 0.2 mHz/rtHz. The loop does not have enough gain to reach these levels.

Quote:

PDH discriminant (40 uV/Hz, see this elog) 

Attachment 1: cavityFlashes.pdf
cavityFlashes.pdf
Attachment 2: modelPDH.pdf
modelPDH.pdf
  14526   Tue Apr 9 00:18:19 2019 gautamUpdateALSEX Green PDH error monitor calibrated

wrong assumption - i checked the gain just now, it is G=10, and is running in the "low-noise" mode, so can only drive 4V. fixed elog, filter.

Note: While working at EX, I saw frequent saturations (red led blinking) on the SR560. Looking a the error mon signal on a scope, it had a pk-to-pk amplitude of ~200mV going into the SR560. Assuming the free-swinging cavity length changes by ~1 um at 1 Hz, the green frequency changes by ~15 MHz, which according to the PDH discriminant calibration of 40 uV/Hz should only make a 60 mV pk to pk signal. So perhaps the cavity length is changing by 4x as much, plausible during daytime with me stomping around the chamber I guess.. My point is that if the SR560 get's saturated (i.e. input > 13000 cts), the DQ-ed spectrum isn't trustworthy anymore. Should hook this up to some proper whitening electronics

Quote:

G=10 or G=100?

  14525   Tue Apr 9 00:16:22 2019 ranaUpdateALSEX Green PDH error monitor calibrated

G=10 or G=100?

  14524   Mon Apr 8 23:52:09 2019 gautamUpdateALSEX Green PDH error monitor calibrated

Some time ago, I had done an actuator calibration of ITMX. This suspension hasn't been victim to the recent spate of suspension problems, so I can believe that the results of those measurement are still valid. So I decided to calibrate the in-loop error signal of the EX green PDH loop, which is recorded via an SR560, G=10, by driving a line in ITMY position (thereby modulating the X arm cavity length) while the EX green frequency was locked to the arm cavity length. Knowing the amount I'm modulating the cavity length by (500 cts amplitude sine wave at 33.14159 Hz using awggui, translating to ~17.2 kHz amplitude in green frequency), I demodulated the response in C1:ALS-X_ERR_MON_OUT_DQ channel. At this frequency of ~33 Hz, the servo gain should be large, and so the green laser frequency should track the cavity length nearly perfectly (with transfer function 1/(1+L), where L is the OLG).

The response had amplitude 5.68 +/- 0.10 cts, see Attachment #1. There was a sneaky gain of 0.86 in the filter module, which I saw no reason to keep at this strange value, and so updated to 1, correcting the demodulated response to 6.6 cts. After accounting for this adjustment, the x10 gain of the SR560, and the loop suppression, I put a "cts2Hz" filter in (Attachment #2). I had to guess a value for the OLG at 33 Hz in order to account for the in-loop suppression. So I measured the OLTF using the usual IN1/IN2 method (Attachment #3), and then used a LISO model of the electronics, along with guesses of the cavity pole (18.5 kHz), low-pass filter poles (4x real poles at 70 kHz), PZT actuator gain (1.7 MHz/V) and PDH discriminant (40 uV/Hz, see this elog) to construct a model OLTF. Then I fudged the overall gain to get the model to line up with the measurement between 1-10kHz. Per this model, I should have ~75dB of gain at ~33Hz, so the tracking error to my cavity length modulation should be ~3.05 Hz. Lines up pretty well with the measured value of 4.7 Hz considering the number of guessed parameters. The measured OLG tapers off towards low frequency probably because the increased loop suppression drives one of the measured inputs on the SR785 into the instrument noise floor.

The final calibration number is 7.1 Hz/ct, though the error on this number is large ~30%. Note that these "Hertz" are green frequency changes - so the change to the IR frequency will be half.

Attachment #4 shows the error signal in various conditions, labelled in the legend. Interpretations to follow.

Attachment 1: errMonCalib.pdf
errMonCalib.pdf
Attachment 2: errMon.png
errMon.png
Attachment 3: OLTF.pdf
OLTF.pdf
Attachment 4: EX_frequencyNoises.pdf
EX_frequencyNoises.pdf
  14523   Mon Apr 8 18:28:25 2019 gautamUpdateALSEX Green PDH checkout

I worked on characterizing the green PDH setup at EX, as part of the ALS noise budgeting process. Summary of my findings:

  1. Green doubling efficiency is ~ 1.5 %/W (3mW of green for 450mW of IR). This is ~half of what was measured on the PSL table. There are probably large errors associated with power measurement with the Ophir power meter, but still, seems like a big mismatch.
  2. The green REFL photodiode is a Thorlabs PDA36A
    • It is being run on 30 dB gain setting, corresponding to a transimpedance of 47.5 kohm into high impedance loads. However, the PD bandwidth for this gain setting is 260 kHz according to the manual, whereas the PDH modulation sidebands on the green light are at twice the modulation frequency, i.e. ~560 kHz, so this is not ideal.
    • There was ~250 uW of green light incident on this photodiode, as measured with the Ophir power meter.
    • The DC voltage level was measured to be ~2.7 V on a scope (High-Z), which works out to ~280 uW of power, so the measurements are consistent.
    • When the cavity is locked, there is about 25% of this light incident on the PD, giving a shot noise level of ~25 nV/rtHz. The dark noise level is a little higher, at 40nV/rtHz.
    • Beam centering on the PD looked pretty good to the eye (it is a large-ish active area, ~3mmx3mm).
    • The beam does not look Gaussian at all - there are some kind of fringes visible in the vertical direction in a kind of halo around the main cavity reflection. Not sure what the noise implications of this are. I tried to capture this in a photo, see Attachment #1. Should an Iris/aperture be used to cut out some of this junk light before the reflection photodiode?
  3. The in-going beam was getting clipped on the Faraday Isolator aperture (it was low in pitch).
    • I fixed this by adjusting the upstream steering, and then moving the two PZT mounted green steering mirrors to recover good alignment to the X arm cavity.
    • GTRX level of ~0.5 was recovered.
  4. To estimate the mode-matching of the input beam to the cavity axis, I looked at the reflected light with the cavity locked, and with just the prompt reflection from the ETM:
    • DC light level on the reflection photodiode was monitored using the High-Z input o'scope.
    • Measured numbers are Plocked ~ 660 mV, Pmisaligned ~ 2.6V, giving a ratio of 0.253.
    • While locked, there was a ~ 10 Hz periodic variation in the DC light level on the green REFL photodiode - not sure what was causing this modulation.
    • However, this is inconsistent with a calculation, see Attachment #2. I assumed modulation depth of 90 mrad, round-trip loss of 100 ppm, and Titm = 1.094%, Tetm = 4.579%, numbers I pulled from the core-optics wiki page.
    • Not sure what effect I've missed out on here - to get the model to match the measurement, I have to either assume a higher cavity finesse, or a much higher round-trip loss (5000ppm), both of which seem implausible.

The main motivation was to get the residual frequency noise of the EX laser when locked to the X arm cavity - but I'll need the V/Hz PDH discriminant to convert the in-loop error signal to frequency units. The idea was to look at the PDH error signal on a scope and match up the horn-to-horn voltage with a model to back out said discriminant, but I'll have to double check my model for errors now given the large mismatch I observe in reflected power.

Attachment 1: IMG_7393.JPG
IMG_7393.JPG
Attachment 2: greenModeMatch.pdf
greenModeMatch.pdf
  14522   Mon Apr 8 11:53:17 2019 gautamUpdateCDSc1oaf needs debugging

I tried restarting c1oaf this weekend to see if turning on the MC length FF would affect the ALS noise performance. I burtrestored the filter settings from March 2016. However, I noticed several possible anomalies, which need debugging. I am not turning the model off because of the possibility of having to reboot all the vertex FEs, but this model is totally unusable right now.

  1. Attachment #1 - the vertex seismometer input produces 1e+20 cts at the output of the feedforward filter. Attachment #2 shows the shape of the feedforward filters - doesn't explain the saturation. Since this is a feedforward loop, a runaway loop can't be the explanation either.
  2. The MC length feedforward control signal is supposed to only go to MC2 - but MC1 and MC3 coil outputs were saturated when I enabled the feedforward.
Attachment 1: c1oaf_sat.png
c1oaf_sat.png
Attachment 2: MCL_FF_TF.pdf
MCL_FF_TF.pdf
  14521   Mon Apr 8 00:04:08 2019 gautamUpdateALSIR ALS noise budget

To start the noise budgeting, I decided to measure the "DFD noise", which is really the quadrature sum of the following terms:

  • ZHL-3A (RF amplifier) noise, NF ~ 6dB per spec (~ 1nV/rtHz)
  • Delay line demod board noise, ~30nV/rtHz [measurement]
  • AA board noise [measurement]
  • ADC noise

According to past characterizations of these noises, the ADC noise level, which is expected to be at the level of a few uV/rtHz, is expected to be the dominant noise source.

The measurement was made by disconnecting the NF 1611 free space photodiode from the input to the RF amplifier on the PSL table, and connecting a Marconi (f_carrier = 40 MHz, signal level=-5dBm) instead. The phase tracked output was then monitored, and the resulting digital spectrum is the red curve in Attachment #1. The blue curve is the ASD of fluctuations of the beatnote between the PSL and EX IR beams, as monitored by the DFD system, with the X arm cavity length locked to the PSL frequency via the LSC servo, and the EX green frequency locked to the X arm cavity length by the analog PDH servo. 

Conclusions:

Assuming the Marconi frquency noise is lower than the ones being budgeted:

  1. the measured frequency noise is above the DFD noise - this needs to be budgeted.
  2. The DFD noise level is consistent with a frequency discriminant of 15 uV/rtHz and an ADC noise level of 3 uV/rtHz at high frequencies.

Next noises to budget:

  1. In-loop X arm length noise
  2. In-lop EX laser frequency noise
Attachment 1: DFDnoise.pdf
DFDnoise.pdf
  14520   Sat Apr 6 02:07:40 2019 AnjaliUpdateFrequency noise measurementFrequency noise measurement of 1 micron source
  • The alignment of the output beam from the delayed path of MZI to the photodetector was disturbed when we did the polarisation characterisation yesterday. So, today we tried to align the output beam from the delayed path of MZI to the detector .
  • We then observed the beat output from the detector on oscilloscope.We initialy observed a dc shift . We then applied a frequency modulation on the input laser and observed the output on oscilloscope. We expected to see variation in output frequency in accordance with variation of input frequency modulation. But we didnt observe this and we were not really getting the interference pattern. 
  • We tried to make the alignment better. With a better alignment, we could see the interference pattern. We also observed that the output frequency was varying in accordance with variation in the input frequency modulation. We would expect a better result with proper mode matching of the two beams on the photodetector.
  14519   Fri Apr 5 11:49:30 2019 gautamUpdateALSPSL + X green beat recovery

Since we haven't been using it, the PID control was not enabled on the doubling oven on the PSL table (it is disabled after every power outage event in the lab). I re-enabled it just now. The setpoint according to the label on the TC200 controller is 36.9 C. The PID paramaters were P=250, I=200, D=40. These are not very good as the overshoot when I turned the control on was 44 C, seems too large. The settling time is also too long, after 10 minutes, the crystal temperature as reported by the TC200 front panel is still oscillating. I can't find anything in the elog about what the nominal PID parameter values were. The X end PID seems much better behaved so I decided to try the same PID gains as is implemented there, P=250, I=60, D=25.

With the Ophir power meter, I measured 60mW of IR light going into the doubling oven, 110uW green light coming out, for a conversion efficiency of 2.7%/W, seems pretty great.

Next, I went to EX and tweaked the steering mirror alignment - I wasn't able to improve the transmission significantly using the PZT sliders on the EPICS screen, and the dither alignment servo isn't working. It required quite a substantial common mode yaw shift of the PZT mirrors to make GTRX ~ 0.5. 

Quote:

I plan to recover the green beat note as well and digitize it using the second available DFD channel (eventually for the Y arm) - then we can simultaneously compare the the green and IR performance (though they will have different noise floors as there is less green light on the green beat PDs, and I think lower transimpedance too).

  14518   Fri Apr 5 11:40:57 2019 AnjaliUpdateFrequency noise measurementFrequency noise measurement of 1 micron source
  • Attachment #1 shows the present experimental setup. The photodiode is now replaced with PDA255. The farther end of the fiber (output of the delayed arm) is coupled through a collimator and aligned such that the beam from the delayed path fall on the detector along with the undelayed path of MZI. We tried to measure the frequency noise of the laser with this setup, but we didn’t get anything sensible.
  • One of the main draw backs of the measurement was the polarisation was not aligned properly in the setup. So, then the next step was to identify the polarisation at different locations in the beam path and to maximise the polarisation to either S or P component.

  • So, we introduced HWP at the input beam path after isolator as shown in attachment #1. Also, the polarisation was tested at positions P1, P2, P3, and P4 shown in attachment #1 by placing a polarisation beam splitter at these locations and then by observing the transmitted (P component) and reflected light (S component) using power meter.

  • The observations at different locations are as the follows

Position Input power (mW) P component (mW) S component (mW)
P1 279 145 123
P2 255 113 137
P3 129 67 58
P4 124 66 53

 

  • These observations show that the P and S components are almost equal, and this is not a good polarisation arrangement. At this point, we also had to check whether the incoming beam is linearly polarised or not.

  • To test the same, the PBS was placed at position P1 and the P and S components were observed with power meter as the HWP is rotated.Attachment # 2 shows the results of the same, that is the variation in P and S component as the HWP is rotated.

  • This result clearly shows that the input beam is linearly polarised. The HWP was then adjusted such that the P component is maximum and coupled to the MZI. With this orientation of HWP, the polarisation observed at different positions P1, P2, P3, and P4 are as follows.

Position Input (mW) P component (mW) S component (mW)
P1 283 276 5
P2 248 228 7
P3 126 121 2
P4 128 117 1
  • This shows that the polarisation is linearly polarised as well as it is oriented along the P direction (parallel to the optical table).

  • We have the polarisation maintaining fiber (PM 980) as the delay fiber. The polarisation of the light as it propagates through a PM fiber depends on how well the input beam is coupled to the axis (slow or fast) of the fiber. So, the next task was to couple the light to one of the axes of the fiber.

  • The alignment key on the fiber is a good indication of the axis of the fiber. In our case, the alignment key lines up with the slow axis of the fiber. We decided to couple the light to the fast axis of the fiber. Since the incoming beam is P polarised, the output fiber coupler was  aligned such that the fast axis is parallel to optical table as possible.

  • A PBS was then introduced after the fiber output collimator . There is a HWP (marked as HWP2 in attachment 1) in front of the input coupler of the fiber as well. This HWP was then rotated and observed the P and S component from the PBS that is now placed after the output coupler with a power meter.The idea was , when the light is coupled to the fast axis of the fiber, we will see the maximum at the P componet at the output

  • Attachment # 3 shows the observation. 

  • In this way I tried to find the orientation of the HWP2 such that the P component is maximum at the output. But I was not succeeded in this method and observed that the output was fluctuating when the fiber was disturbed. One  doubt we had was whether the fiber is PM or not . Thus we checked the fiber end with fiber microscope and confirmed that it is PM fiber. 

  • Thus, we modifed the setup as shown in attachement # 4.The photodetector (PDA55) was monitoring the S component and the output of the detector was observed on an oscilloscope. We rotated the HWP2 such that the S component is almost minimum. At the same time, we were disturbing the fiber and was observing whether the output is fluctuating. The HWP2 angle was tweaked around the minimum of S component and observed the output with disturbing the fiber. This way we found the orientation of HWP2  such that the light is coupled to the fast axis of the fiber and the output was not fluctuating while we disturb the fiber. We tested it  by heating the fiber with a heat gun as well and confirmed that the output is not fluctuating and thus the light is coupled to the fast axis of the fiber.

Attachment 1: Modified_experimental_setup.JPG
Modified_experimental_setup.JPG
Attachment 2: Checking_polarisation.pdf
Checking_polarisation.pdf
Attachment 3: Checking_the_polarisation_alignment_of_the_delay_fiber.pdf
Checking_the_polarisation_alignment_of_the_delay_fiber.pdf
Attachment 4: Setup_to_test_the_polarisation_alignment_of_delay_fiber.JPG
Setup_to_test_the_polarisation_alignment_of_delay_fiber.JPG
  14517   Fri Apr 5 01:10:18 2019 gautamUpdateVACTP3 forepump is also noisy

Is this one close to failure as well?

  14516   Fri Apr 5 00:33:58 2019 gautamUpdateALSPromising IR ALS noise

Summary:

I set up a free-space beat on theNW side of the PSL table between the IR beam from the PSL and from EX, the latter brought to the PSL table via ~40m fiber. Initial measurements suggest very good performance, although further tests are required to be sure. Specifically, the noise below 10 Hz seems much improved.

Details:

Attachment #1 shows the optical setup. 

  • I used two identical Thorlabs F220APC collimators to couple the light back into free space, reasoning that the mode-matching would be easiest this way.
  • Only 1 spare K6Xs collimator mount was available (this has the locking nut on the rotational DoF), so I used a K6X for the other mount. The fast axis of the Panda fibers were aligned as best as possible to p-polarization on the table by using the fact that the connector key is aligned to the slow axis.
  • I cut the power coupled into the PSL fiber from ~2.6mW to ~880uW (using a HWP + PBS combo before the input coupling to the fiber) to match the power from EX.
  • The expected signal level from these powers and the NF1611 transimpedance of 700 V/A is ~320 mVpp. After alignment tweaking, I measured ~310mVpp (~ -5dBm) into a 50 ohm input on a scope, so the mode-matching which means the polarization matching and mode overlap between the PSL and EX beams are nearly optimal.
  • To pipe the signal to the delay line electronics, I decided to use the ZHL-3A (G=27dB, 1dB compression at 29.5dBm per spec), so the signal level at the DFD rack was expected (and confirmed via 50 ohm input on o'scope) to be ~19dBm.
  • This is a lot of signal - after the insertion loss of the power splitter, there would still be ~15dBm of signal going to the (nominally 10dBm) LO input of the demod board. This path has a Teledyne AP1053 at the input, which has 10dB gain and 1dBm compression at 26dBm per spec. To give a bit of headroom, I opted on the hacky solution of inserting an attenuator (5dB) in this path - a better solution needs to be implemented.
  • The differential outputs of the demod board go to the CDS system via an AA board (there is no analog whitening).

Yehonathan came by today so I had to re-align the arms and recover POX/POY locking. This alllowed me to lock the X arm length to the PSL frequency, and lock the EX green laser to the X arm length. GTRX was ~0.36, whereas I know it can be as high as 0.5, so there is definitely room to improve the EX frequency noise suppression.

Attachment #2 shows the ALS out-of-loop noise for the PSL+X combo. The main improvements compared to this time last year are electronic. 

  • The failed experiment of making custom I/F amplifier was abandoned and Rich Abbott's original design was reverted to. 
  • New power splitter was installed with 3dB less insertion loss.
  • According to the RF path level monitor, the signal level at the RF input to the demod board is ~10dBm. Per my earlier characterization, this will give us the pretty beefy frequency discriminant of ~15uV/Hz.
  • I estimate the frequency noise of the detection electronics + ADC noise now translate to 1/3 the frequency noise compared to the old system. With some analog whitening, this can be made even better, the electronics noise of the DFD electronics (~50nV/rtHz) is estimated to be <10mHz/rtHz equivalent frequency noise. 
  • Note that the calibration from phase-tracker-servo to units of Hz (~14 kHz / degree) was not changed in the digital system - this should only be a property of the delay line length, and hence, should not have changed as a result of the various electronics changes to the demod board and other electronics.

Next steps:

  • Improve pointing of green beam into X arm cavity.
  • I plan to recover the green beat note as well and digitize it using the second available DFD channel (eventually for the Y arm) - then we can simultaneously compare the the green and IR performance (though they will have different noise floors as there is less green light on the green beat PDs, and I think lower transimpedance too).
Quote:

Mix the beams in free space. We have the beam coming from EX to the PSL table, so once we mix the two beams, we can use either a fiber or free-space PD to read out the beatnote. 

  • This approach means we lose some of the advantages of the fiber based setup (e.g. frequent alignment of the free-space MM of the two interfering beams may be required).
  • Potentially increases sensitivity to jitter noise at the free-space/fiber coupling points
Attachment 1: IMG_7388.JPG
IMG_7388.JPG
Attachment 2: freeSpace_IR_beat.pdf
freeSpace_IR_beat.pdf
  14515   Wed Apr 3 18:35:54 2019 gautamUpdateVACPSL shutter re-opened

PSL shutter was re-opened at 6pm local time. IMC was locked. As of 10pm, the main volume pressure is already back down to the 8e-6 level.

  14514   Wed Apr 3 16:17:17 2019 JonUpdateVACTP2 forepump replaced

I can't explain the mechanical switching sound Gautam reported. The relay controlling power to the TP2 forepump is housed in the main AC relay box under the arm tube, not in the Acromag chassis, so it can't be from that. I've cycled through the pumpdown sequence several times and can't reproduce the effect. The Acromag switches for TP2 still work fine.

In any case, I've made modifications to the vacuum interlocks that will help with two of the issues:

  1. For the "AC power loss" over-triggering: New logic added requiring the UPS to be out of the "on line power, battery OK" state for ~5 seconds before tripping the interlock. This will prevent electrical transients from triggering an emergency shutdown, as seems to be the case here (the UPS briefly isolates the load to battery during such events).
  2. PSL interlocking: New logic added which directly sets C1:AUX-PSL_ShutterRqst --> 0 (closes the PSL shutter) when the main volume pressure is 3 mtorr-500 torr. Previously there was a channel exposed for this interlock (C1:Vac-interlock_high_voltage), but c1aux was not actually monitoring it. Following the convention of every vac interlock, after the PSL shutter has been closed, it has to be manually reopened. Once the pressure is out of this range, the vac system will stop blocking the shutter from reopening, but it will not perform the reopen action itself. gautam: a separate interlock logic needs to be implemented on c1aux (the shutter machine) that only permits the shutter to be opened if the Vac pressure range is okay. The SUS watchdog style AND logic in the EPICS database file should work just fine.

After finishing this vac work, I began a new pumpdown at ~4:30pm. The pressure fell quickly and has already reached ~1e-5 torr. TP2 current and temp look fine.

Quote:

However, when opening V4 (the foreline of TP1 pumped by TP2), I heard a loud repeated click track (~5Hz) from the electronics rack. Shortly after, the interlocks shut down all the TPs again, citing "AC power loss". Something is not right, I leave it to Jon and Chub to investigate.

Attachment 1: IMG_3180.jpg
IMG_3180.jpg
  14513   Wed Apr 3 12:32:33 2019 KojiUpdateALSNote about new fiber couplers

Andrew seems to have an integrated solution of PBS+HWP in a singe mount. Or, I wonder if we should use HWP/QWP before the coupler. I am interested in a general solution for this problem in my OMC setup too.

  14512   Wed Apr 3 10:42:36 2019 gautamUpdateVACTP2 forepump replaced

Bob and Chub concluded that the drypump that serves as TP2's forepump had failed. Steve had told me the whereabouts of a spare Agilent IDP-7. This was meant to be a replacement for the TP3 foreline pump when it failed, but we decided to swap it in while diagnosing the failed drypump (which had 2182 hours continuous running according to the hour counter). Sure enough, the spare pump spun up and the TP2fl pressure dropped at a rate consistent with what is expected. I was then able to spin up TP1, TP2 and TP3. 

However, when opening V4 (the foreline of TP1 pumped by TP2), I heard a loud repeated click track (~5Hz) from the electronics rack. Shortly after, the interlocks shut down all the TPs again, citing "AC power loss". Something is not right, I leave it to Jon and Chub to investigate.

  14511   Wed Apr 3 09:07:46 2019 gautamUpdateVACVac failure

Overnight pressure trends don't suggest anything went awry after the initial interlock trip. Some watchdog script that monitors vacuum pressure and closes the PSL shutter in the event of pressure exceeding some threshold needs to be implemented. Another pending task is to make sure that backup disk for c1vac actually is bootable and is a plug-and-play replacement.

Attachment 1: vacFailOvernight.png
vacFailOvernight.png
  14510   Wed Apr 3 09:04:01 2019 gautamUpdateALSNote about new fiber couplers

The new fiber beam splitters we are ordering, PFC-64-2-50-L-P-7-2-FB-0.3W, have the slow axis working and fast axis blocked. The way the light is coupled into the fibers right now is done to maximize the amount of light into the fast axis. So we will have to do a 90deg rotation if we use that part. Probably the easiest thing to do is to put a HWP immediately before the free-space-to-fiber collimator.

Update 6pm: They have an "SB" version of the part with the slow axis blocked and fast axis enabled, same price, so I'll ask Chub to get it.

  14509   Tue Apr 2 18:40:01 2019 gautamUpdateVACVac failure

While glancing at my Vacuum striptool, I noticed that the IFO pressure is 2e-4 torr. There was an "AC power loss" reported by C1Vac about 4 hours (14:07 local time) ago. We are investigating. I closed the PSL shutter.


Jon and I investigated at the vacuum rack. The UPS was reporting a normal status ("On Line"). Everything looked normal so we attempted to bring the system back to the nominal state. But TP2 drypump was making a loud rattling noise, and the TP2 foreline pressure was not coming down at a normal rate. We wonder if the TP2 drypump has somehow been damaged - we leave it for Chub to investigate and give a more professional assessment of the situation and what the appropriate course of action is.

The PSL shutter will remain closed overning, and the main volume and annuli are valved off. We spun up TP1 and TP3 and decided to leave them on (but they have negligible load).

Attachment 1: vacFail.png
vacFail.png
  14508   Tue Apr 2 15:02:53 2019 JonUpdateCDSITMY freed

blushI renamed all channels on c1susaux2 from "C1:SUS-..." to "C1:SUS2-..." to avoid contention. When the new system is ready to install, those channel names can be reverted with a quick search-and-replace edit.

Quote:

While doing this work, I noticed several errors corresponding to EPICS channel conflicts. Turns out the c1susaux2 EPICS server was left running, and the MEDM screens (and possibly several scripts) were confused. There has to be some other way of testing the new crate, on an isolated network or something - please do not leave the modbus service running as it potentially interferes with normal IFO operation. For good measure, I stopped the process and shut down the machine since I saw nothing in the elog about any running tests.

  14507   Tue Apr 2 14:53:57 2019 gautamUpdateCDSc1vac added to burt

I deleted references to c1vac1 and c1vac2 (which no longer exist) and added c1vac to the autoburt request file list at /opt/rtcds/caltech/c1/burt/autoburt/requestfilelist

  14506   Mon Apr 1 22:33:00 2019 gautamUpdateCDSITMY freed

While Anjali is working on the 1um MZ setup, the pesky ITMY was liberated from the OSEMs. The "algorithm" :

  • Apply a large (-30000 cts) offset to the side coil using the fast system.
  • Approach the zero of the YAW DoF from -2.00V, PIT from +10V (you'll have to jiggle the offsets until the optic is free swinging, and then step the bias down by 0.1). At this point I had the damping off.
  • Once the PIT bias slider reaches -4V, I engaged all damping loops, and brought the optic to its nominal bias position under damping. 

While doing this work, I noticed several errors corresponding to EPICS channel conflicts. Turns out the c1susaux2 EPICS server was left running, and the MEDM screens (and possibly several scripts) were confused. There has to be some other way of testing the new crate, on an isolated network or something - please do not leave the modbus service running as it potentially interferes with normal IFO operation. For good measure, I stopped the process and shut down the machine since I saw nothing in the elog about any running tests.

Quote:

ITMY became stuck during this process

  14505   Mon Apr 1 12:01:52 2019 JonUpdateCDS 

I brought c1susaux back online this morning for suspension-channel test scripting. It had been dead for some time. I followed the procedure outlined in #12542. ITMY became stuck during this process, which Gautam tells me always happens since the last vacuum access, but ITMX is not stuck.

ELOG V3.1.3-